OpenStudy (anonymous):

an athlete jumps with the speed of 4 m/s at an angle of 23° how long does the athlete stay in the air?

OpenStudy (anonymous):

Okay, this is something I call classic projectile. 1. Method 1: The easy way. We have a result derived from eqn of motion as: \[T = 2u \sin x / g\] Just put the values. 2. Method 2: The hard way. The vertical displacement is zero. Thus, \[S _{y} = 0 = u _{y}t - 1/2 g t ^{2}\] Use \[u _{y} = u \sin x\]. Put values and get the answer. The answer is 0.64. Check it though, I mess up real bad at calculations. :)

rebeccaxhawaii (rebeccaxhawaii):

\(\Huge\boxed{\boxed{\boxed{\huge\mathbb{WELCOME~TO~OPENSTUDY} }}}\) Hey i have made this to hopefully answer any questions you may have. But if it doesn't feel free to pm me c: http://openstudy.com/users/rebeccaxhawaii#/updates/56f5205ce4b07a8c82287f75 Good Luck with you studies! \(\Huge\heartsuit\)